A rigid cylinder of volume 0.028 m3 contains steam at 80 bar and 350°C.

The cylinder is cooled until the pressure is 50 bar. Calculate :

(i) The state of steam after cooling ;

(ii) The amount of heat rejected by the steam.


1
Expert's answer
2021-08-31T10:49:44-0400

Gives

Volume (v)=0.028m3

"P_1=80" bar

T=315°C


"P_2=50 bar"

Mass="\frac{0.028}{0.02995}=0.935kg"

Internal energy


"u_1=h_1-p_1V=2987.3-80\times10^5\times\frac{0.02995}{1000}"

"u_1=2747.7Kj\/kg"

State of steam after cooling


"x_2=\frac{v_2}{v_{g2}}=\frac{0.02995}{0.0394}=0.76"

amount of heat rejected by the steam


"u_2=(1-x_2)u_{f2}+x_2u_{g2}\\u_2=(1-0.76)\times1149+0.76\times2597=2249.48Kj/kg"